recommissioning telescope after painting project
last updated 15dec07
The painting project wound down at the end of
November. We started recommissioning the telescope the 3rd of dec. The
- 04dec07: emergency stop comes on.
- emergency stop problem was a broken cable for the control room
- The control room emergency stop goes from the control room to
box at the rotary joint. This cable was ok. It then goes along a
cable tray under the az walkway over to the vertex shelter. The cable
was broken in the tray.
- did az
swings with tilt sensors to try and level the platform.
- added the following offsets to the tiedown ref positions from
the swing measurements.
- [.186,.272,-.458] (td12,td4,td8)
- az stops. motor 42 failure. replaced amp..
- When using program track mode az would move to lower az when
commanded to sit still. This has occurred before.
- replaced plc,lcu to get rid of program tracking problem.
- Got some more motor failures, a few brake failures.
- checked the brakes on 4 of the az motors. they were clean
- check the resolver cables that went to each az motor (at the
motor end). They were dirty and they all had to be cleaned. One of the
pins on motor 42 resolver cable (sinHi) looked like it was pushed in a
bit. This could have been causing the amp 42 failures (motor over
- more az
swings with tilt sensors to see how well the 06dec07 td adjustments
- The 1 az term for pitch and roll were now about .002 degrees.
- using sbn tracked 3
sources (.ps) (.pdf)
- rms pointing error was about 5 Asecs. processing: x101/071207/chkpnt.pro
- operator computer dies about 1am. lost scramnet. B0732+332 had
no az,za positions.
- for now we've but a jumper in the scramnet fiber to bypass
- radar experiment AM. got cw and 1 image before pfs failed and
cbr disc trouble.
- maintenance lower the tertiary control cabinet that was in the
dome. Added 800 lbs of lead for the new tertiary control box that will
eventually go there.
- using sbn tracked 7
sources (.ps) (.pdf)
14:00 to 5am next day to check pointing. Had hard rain off and
- rms pointing error: 5.2 az, 6.6 az. Looks like we have a ramp
in az/za error vs za. Probably because alfa is not up yet.
- around 5am pointing check halted with bending error.
- the jump was on the ch encoder. Went up about 6am and rotated
it back. It jumped multiple teeth. The rack gear looked good at this
- Started radar run.
- Az bending again. Ch1 encoder again. az was at 327
degrees (dome side).
- went up and found that the encoder was not engaged in the
rack gear. I pushed it forward and it went in. Rack gear teeth looked
good at this position. This spot is where the rack gear moves out by
about 3/4 inches.
- Maintenance put some silicone grease on the spring mechanism.
Need to take it apart and have a closer look. Should do this at 327 so
we can push it all the way forward to see if it sticks.
- raised the ramp for alfa.
the nuts on the tiedowns so that they can be put at the
same extension (there had been a 3.5 inch difference).
- moved nuts for td8 up 3.25 inches. moved nuts for td 4 down
.5 inches. td 12 was not changed.
- doing az swings dome at 15 degrees to see if encoder fails
- Can't use tilt sensors since alfa ramp is on the rotary floor
preventing rotation to 75Deg for the tilt sensor usage.
- alfa was lifted and installed.
- Found that az enc2 cover was stopping the encoder mount from
- Horn survey will be done, but we probably won't use the results
to move the horns.
- The reference targets on the secondary have changed a little
- We have to survey the tertiary (videogrammetry and install
the tertiary target paddles) so we know the tertiary location in this
"new" system before we reshoot the horns.
- Checking the tiedown positions after the td nuts were moved.
- Az bending error on last
set of az swings. az enc2 (ch side) came off out of the rack gear again.
- Raised another compressor.
- wait for the rain to pass.
- Az motor 81 has toruqes that are higher than the other motors. (torques for motor 81 are
too high). Opened the motor and found that the brake had worn.
Replaced the brake pad.
- Pcu portable control units not working in any of the 3
- More bending errors on encoder 2. Dismantled the moveable base
that keeps the encoder in contact with the rack gear. Found grit in the
ball bearings of the slide mechanism (from the sand blasting). Switched
to the auxiliary mount on the left of the alignment pointer. Ran for 24
hours pointing without any more problems. The encoder mounts need to be
dismantled and cleaned.
- Greased the azimuth rails.
- After greasing we got motor overspeed when in preset mode.
- The wheels were spinning on the grease.
- The problem occurred in preset mode. This mode uses the PI
- Switched to open loop rate mode and the telescope moved
without a problem
- Did a complete 360 degree spin in rate mode to clean off
the grease. Came back in preset (closed loop) mode and things worked ok.
- Checked the range of the encode mount as we spun around.
- Came within .25 inches of the outside stop.
- Measured the brake release currents in 4 motors.
- At the cable where it enters the vertex shelter. saw 1 amp
which is expected.
- At the contacts for the brake relays (behind the lcu). There
was an amp difference +1.4, -.6 for two sets of brakes. This needs to
be checked later.
- Started the pointing model at 17:00. next morning found that
the tiedowns were not tracking the platform height. Looking at the
pointing erros, the data can't be used for the model.
- replaced card in pcu. Tested on telescope and it now works.
- tracked sources for the pointing model
- az drive stopped 1am.. motor 12 not ready. At amplifier showed
overspeed. spent most of the day working on it.
- checked the resolver cables vertex to motor.
- continuity ok, no shorts to sheild, shield intact.
- replaced amplifier, error persisted.
- replaced the motor, the error persisted.
- moved amplifier to motor 11 and overspeed went with the
- Found that some azimuth amplifiers had compensation cards
- The compensation cards set the gain and torque for the
amps. If we were using the wrong comp cards then we could get
overspeed. The problem might be worse now since we've greased the track.
- Comparing the spare compensation cards:
- The gr,ch should have the same resistors.
- The az resistors are different from the gr,ch.
- The spare az cards do not all have the same resistors
- the spare gr cards to not all have the same resistors.
- The vertex manual has a single set of resistor values for
the compensation cards. It does not say anything about gr,az being
- Jorge thinks there is some documentation somewhere that
specifies the correct resistors.
- Removed an amplifier from ch to test on dome.. didn't work.
Ch offline till we reinstall the amplifier.
- Measured the speed output of amp 12 and found it higher than
all the others of the same kind (outer wheel). It had a voltage of 2.4
volts while the others were 1.7. This is probaby part of the overspeed
problem since all the outer wheels have to move at the same speed.
- To do the pointing runs we switched to auxilliary mode. This
drives with 1/2 the az motors (41,42,81,82).
- load cell 2 on td 4 is too high by 17 kips. This has caused
then tension on the block to go over 100Kips (when it was really on 83
kips). The operator stopped the tiedown tracking and sent the tiedowns
to a fixed position to make the tension go down. This lowered the
and wrecked the pointing runs for weds and a few thurs nite. Operators
now know to not stop the td tracking. At some point we should
recalibrate the load cell.
- Tracked sources for the pointing model.
- tracked sources for the pointing model with az motors in aux
- To do:
- figure out the correct resistors for the amp compensation
boards. See if this fixes the overspeed problem.
- realign the distomats
- recalibrate loadcells td 4